# On the contraction properties of Sinkhorn semigroups

### v2

@@ 2.1 Sinkhorn bridges
$U_{2n+1}$ $=$ $~U_{2n}~:=U+\log{{\cal Q}(e^{-V_{2n}})}$
$V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\cal R}(e^{-U_{2n}})}.$ (18)
When $n\to\infty$, the sequence of distributions ${\cal P}_{2n}$ converges towards the Schrödinger bridge (4) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t. the reference measure ${\cal P}_{0}$) while the sequence ${\cal P}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t. the reference measure ${\cal P}_{1}$).
Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (4) can also be computed from any Sinkhorn bridge reference distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
$\mathbb{P}=\operatorname*{arg\,min}_{{\cal P}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\cal P}~|~{\cal P}_{2n}).$ (19)
@@
${\cal S}_{2n}(x,dy)=\frac{{\cal Q}(x,dy)e^{-V_{2n}(y)}}{{\cal Q}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\cal S}_{2n+1}(y,dx)=\frac{{\cal R}(y,dx)~e^{-U_{2n+1}(x)}}{{\cal R}(e^{-U_{2n+1}})(y)},$ (20)
while Proposition 6.3 in [adm-24] yields the formulae
$V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$ (21)
for the potentials $(U_{n},V_{n})$, $n\geq 0$, expressed in terms of the distribution flow of the internal states of the Gibbs-loop process.
2.2 Sinkhorn semigroups
For any $n\geq 0$ we have the fixed point equations
$\lambda_{U}{\cal S}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\cal S}^{\circ}_{2(n+1)}=\nu_{V}$ (22)
with the forward-backward Gibbs-type Markov transitions
${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}.$
The above fixed point equations yield, for any $n\geq 1$, the time inhomogeneous Markov evolution equations
$\pi_{2n}=\pi_{2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\cal S}^{\circ}_{2n+1}.$ (23)
For any $0\leq p\leq n$ we have the semigroup equations
$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
The Sinkhorn transitions ${\cal S}^{\circ}_{2p,2n}$ and ${\cal S}^{\circ}_{2p+1,2n+1}$ are defined inductively by the time inhomogeneous semigroup equations
${\cal S}^{\circ}_{2p,2n}={\cal S}^{\circ}_{2p,2(n-1)}{\cal S}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\cal S}^{\circ}_{2p+1,2n+1}={\cal S}^{\circ}_{2p+1,2n-1}{\cal S}^{\circ}_{2n+1},$
with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator. The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (23) sharing the same invariant measures (22) at each individual step.
@@ 2.2 Sinkhorn semigroups
At every time step $n\geq 0$, the conjugate formulae (16) also show that ${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure
${\cal P}_{2n}(d(x,y))=\lambda_{U}(dx)~{\cal S}_{2n}(x,dy)=\pi_{2n}(dy)~{\cal S}_{2n+1}(y,dx)$
For $n=0$, the two-blocks Gibbs sampler with the forward/backward transition ${\cal S}^{\circ}_{1}:={\cal S}_{0}{\cal S}_{1}$ and target product measure ${\cal P}_{0}=\lambda_{U}\times{\cal S}_{0}$ can be interpreted as a sampling technique for $\lambda_{U}$ based on data augmentation or auxiliary variables (a.k.a. proximal sampling), see for instance [damlen1999gibbs, rendell2020global, vono2020asymptotically, vono2019split] and references therein. For a more detailed discussion on proximal samplers we also refer to the series of articles [mou, chewi-chen, chewi-phd, jiaming, jiaojiao, lee, guan]. Similarly, ${\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure ${\cal P}_{2n+1}$.
2.3 Some stability theorems
In contrast with Kantorovich criteria or related Wasserstein-type semi-distances, even when $\mathbb{X}=\mathbb{Y}$ and the marginals coincide, $\lambda_{U}=\nu_{V}$, the identity-map deterministic coupling has infinite entropy. That is, for any reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ of the form (3) we have
${\cal P}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=+\infty.$
This fact, together with (5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\cal P}_{2n}$ strongly depend on the form of the reference transition $\cal Q$} and the transition potential function $W$ discussed in (2).
Markov transitions of the form (2) with a bounded potential $W$ typically arise on finite or compact spaces. For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos]. The convergence analysis of Sinkhorn bridges in this context is often developed using Hilbert Hilbert projective metrics [chen, deligiannidis, franklin, marino].
In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (13).
@@ 2.3 Some stability theorems
$D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$ (24)
The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$. Similar estimates hold for $(\lambda_{U},\pi_{2n+1})$. There also exists $c>0$ such that for any $n\geq 0$ we have
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{n}.$ (25)
The estimates (24) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1. A more refined version of the entropy estimate (25) in terms of Dobrushin contraction coeffcients is provided in Corollary 4.2
@@
${\cal Q}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$ (26)
@@
${\cal U}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\cal V}_{\delta}:=\delta V-W_{U}$ (27)
@@
$W_{U}(y):=\int~\lambda_{U}(dx)~W(x,y)\quad\text{and}\quad W^{V}(x):=\int~\nu_{V}(dy)~W(x,y).$ (28)
@@
Assume that $W$ is locally bounded and bounded below by a real number and $(U,V)$ are locally bounded with compact sub-level sets. In addition, there exists some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta/2},{\cal V}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$. Then, there is some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ s.t. for any $n\geq n_{0}$ we have
@@
When $\psi$ satisfies the lower bound estimate (9) for some semi-metric $\varrho$ we also have the exponential Wasserstein estimates
$\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$ (30)
@@
The estimates (29) are a consequence of Corollary 4.15 and the Wasserstein formulation (7) of weighted norms. The estimates (30) are a consequence of (10) and (29). More refined strict contraction estimates are provided in Corollary 4.14 and Corollary 4.16.
2.4 Illustrations and comments
@@
In Section 5, the stability theorems presented in Section 2.3 and Section 4 are illustrated in a variety of situations ranging from polynomial growth potentials (Section 5.1) and heavy tailed marginals on general normed spaces (Section 5.2), to more sophisticated boundary state space models (Section 5.3), semi-compact models including semi-discrete target spaces (Section 5.7) as well as finite mixture of statistical models (Section 5.9). Section 5.8 is dedicated to explicit descriptions and equivalent regularity conditions in the context of linear Gaussian models.
When $W$ is bounded, it does not appear useful to consider weighted total variation norms. In this context, the stability of Sinkhorn semigroups w.r.t. Hilbert’s metric is well known [borwein, chen, deligiannidis, franklin] and already yields exponential decays w.r.t. the total variation norm. On bounded subspaces $\mathbb{X}\subset\mathbb{R}^{d}$ the total variation norm also dominates Wasserstein distances w.r.t. the Euclidian norm but in general the target space $\mathbb{X}$ is non complete and then not Polish w.r.t. the Euclidean distance. In this context, it is desirable to find a complete distance on the target space that makes the corresponding Wasserstein distance in distribution space complete [bolley2008separability]. It is easy to construct examples of marginal models on non complete metric spaces. For instance, consider the marginal Beta distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
@@
Applying (24) to the total variation distance (see (14)) we readily check the Wasserstein exponential decays
@@
Note that the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric but it is complete for the metric defined as
@@
for some $c_{\delta}>0$. In this context, the l.h.s. estimate (29) yields the $\varrho_{c,\delta}$-Wasserstein estimates
@@
Theorem 2.3 also applies to Gaussian marginal distributions on $\mathbb{Y}=\mathbb{X}=\mathbb{R}^{d}$ and a transition potential $W$ with at most quadratic growth (cf. Section 5.1). In this context, applying (11) to $(p,q)=(2,1)$ we have the $2$-Wasserstein estimate (30), which implies
@@
Section 5.8 provides an explicit description of the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ in the context of linear-Gaussian models. For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\cal Q$} with variance $\tau>2$ we have
${2}/{\tau}<\delta<1\Longrightarrow{\cal U}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$. }$
A symmetric Gaussian reference transition $\cal Q$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$. In this context, the exponential stability of Sinkhorn bridges w.r.t. the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [del2025stability, chiarini, durmus]. The article [adm-24] also provides Wasserstein exponential decays for any Gaussian marginals and any linear-Gaussian reference transition. For these classes of log-concave-type models, our Lyapunov approach yields new weighted total variation norm estimates for sufficiently large time horizons.
@@
Section 2.3 and Section 5 discuss a series of application models that can be handled using our Lyapunov approach. This section is concerned with some limitations of our approach and related important open research problems.
@@
$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
This rather elementary example is a a slight modification of (26) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
@@
In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1). Nevertheless for uniformly bounded transition potential functions, Theorem 2.2 applies to any class of marginal distributions.
@@
3 Contraction of Markov operators

### v3

@@ 2.1. Sinkhorn bridges
$U_{2n+1}$ $=$ $~U_{2n}~:=U+\log{{\mathcal{Q}}(e^{-V_{2n}})}$
(2.4) $V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\mathcal{R}}(e^{-U_{2n}})}.$
When $n\to\infty$, the sequence of distributions ${\mathcal{P}}_{2n}$ converges towards the Schrödinger bridge (1.4) from $\lambda_{U}$ to $\nu_{V}$ (w.r.t. the reference measure ${\mathcal{P}}_{0}$) while the sequence ${\mathcal{P}}_{2n+1}$ converges towards the Schrödinger bridge from $\nu_{V}$ to $\lambda_{U}$ (w.r.t. the reference measure ${\mathcal{P}}_{1}$).
Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (1.4) can also be computed from any Sinkhorn bridge reference distribution ${\mathcal{P}}_{2n}$, in the sense that for any $n\geq 0$ we have
(2.5) $\mathbb{P}=\operatorname*{arg\,min}_{{\mathcal{P}}\,\in\,\Pi_{U,V}}\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{2n}).$
@@
(2.6) ${\mathcal{S}}_{2n}(x,dy)=\frac{{\mathcal{Q}}(x,dy)e^{-V_{2n}(y)}}{{\mathcal{Q}}(e^{-V_{2n}})(x)}\quad\mbox{and}\quad{\mathcal{S}}_{2n+1}(y,dx)=\frac{{\mathcal{R}}(y,dx)~e^{-U_{2n+1}(x)}}{{\mathcal{R}}(e^{-U_{2n+1}})(y)},$
Moreover, Proposition 6.3 in [adm-24] shows that the potentials can be expressed through the marginal flows as
(2.7) $V_{2n}=V_{0}+\sum_{0\leq p<n}\log\frac{d\pi_{2p}}{d\nu_{V}}\quad\mbox{and}\quad U_{2n}=U_{0}+\sum_{0\leq p<n}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}$
2.2. Sinkhorn semigroups
The Sinkhorn iterations naturally induce a family of time-inhomogeneous Markov semigroups. For each $n\geq 0$ define the forward-backward Gibbs-type transitions
${\mathcal{S}}^{\circ}_{2n+1}:={\mathcal{S}}_{2n}{\mathcal{S}}_{2n+1}\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2(n+1)}:={\mathcal{S}}_{2n+1}{\mathcal{S}}_{2(n+1)}.$
These kernels satisfy the fixed-point relations
(2.8) $\lambda_{U}{\mathcal{S}}^{\circ}_{2n+1}=\lambda_{U}\quad\mbox{\rm and}\quad\nu_{V}{\mathcal{S}}^{\circ}_{2(n+1)}=\nu_{V}$
and yield the marginal evolution equations
(2.9) $\pi_{2n}=\pi_{2(n-1)}{\mathcal{S}}^{\circ}_{2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2n-1}{\mathcal{S}}^{\circ}_{2n+1}.$
For $0\leq p\leq n$, we define the associated inhomogeneous semigroups by composition,
${\mathcal{S}}^{\circ}_{2p,2n}={\mathcal{S}}^{\circ}_{2p,2(n-1)}{\mathcal{S}}^{\circ}_{2n}\quad\mbox{\rm and}\quad{\mathcal{S}}^{\circ}_{2p+1,2n+1}={\mathcal{S}}^{\circ}_{2p+1,2n-1}{\mathcal{S}}^{\circ}_{2n+1},$
with identity operators on the diagonal. This representation highlights that Sinkhorn dynamics belong to the class of time-varying Markov chains sharing common invariant measures at each step.
@@ 2.2. Sinkhorn semigroups
Each kernel ${\mathcal{S}}^{\circ}_{2n+1}:={\mathcal{S}}_{2n}{\mathcal{S}}_{2n+1}$ corresponds to a two-block Gibbs sampler with target distribution
${\mathcal{P}}_{2n}(d(x,y))=\lambda_{U}(dx)~{\mathcal{S}}_{2n}(x,dy)=\pi_{2n}(dy)~{\mathcal{S}}_{2n+1}(y,dx)$
For $n=0$, this recovers classical proximal and data-augmentation samplers; see, e.g., [damlen1999gibbs, rendell2020global, vono2019split, vono2020asymptotically] and related works [chewi-chen, chewi-phd, jiaojiao, guan, lee, jiaming, mou].
2.3. Some stability theorems
In contrast with Kantorovich criteria or related Wasserstein-type semi-distances, even when $\mathbb{X}=\mathbb{Y}$ and the marginals coincide, $\lambda_{U}=\nu_{V}$, the identity-map deterministic coupling has infinite entropy. That is, for any reference measure ${\mathcal{P}}_{0}=(\lambda_{U}\times{\mathcal{Q}})$ of the form (1.3) we have
${\mathcal{P}}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\mathcal{P}}~|~{\mathcal{P}}_{0})=+\infty.$
This fact, together with (1.5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\mathcal{P}}_{2n}$ crucially depend on the form of the reference transition $\mathcal{Q}$} and the transition potential function $W$ discussed in (1.2).
Markov transitions of the form (1.2) with a bounded potential $W$ typically arise on finite or compact spaces. For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (1.2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos]. The convergence analysis of Sinkhorn bridges in this context is often developed using Hilbert Hilbert projective metrics [chen, deligiannidis, franklin, marino].
In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (1.13).
@@ 2.3. Some stability theorems
(2.10) $D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$
The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$. Analogous bounds hold for $(\lambda_{U},\pi_{2n+1})$. Moreover,
(2.11) $\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\rho^{n}.$
The estimates (2.10) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1. A more refined version of the entropy estimate (2.11) in terms of Dobrushin contraction coefficients is provided in Corollary 4.2
@@
(2.12) ${\mathcal{Q}}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$
@@
(2.13) ${\mathcal{U}}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\mathcal{V}}_{\delta}:=\delta V-W_{U}$
@@
(2.14) $W_{U}(y):=\int~\lambda_{U}(dx)~W(x,y)\quad\text{and}\quad W^{V}(x):=\int~\nu_{V}(dy)~W(x,y).$
@@
Assume that $W$ is locally bounded and bounded from below by a real number and $(U,V)$ are locally bounded with compact sub-level sets. In addition, there exists some $\delta\in]0,1[$ such that the functions $({\mathcal{U}}_{\delta/2},{\mathcal{V}}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$. Then, there exist some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ such that for any $n\geq n_{0}$,
@@
When $\psi$ satisfies the lower bound estimate (1.9) for some semi-metric $\varrho$ then we also have the exponential Wasserstein estimates
(2.16) $\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$
@@
The estimates (2.15) are a consequence of Corollary 4.10 and the Wasserstein formulation (1.7) of weighted norms. The estimates (2.16) are a consequence of (1.10) and (2.15). More refined strict contraction estimates are established in Corollary 4.9 and Corollary 4.11.
2.4. Illustrations and comments
@@
Section 5 provides a range of examples illustrating the stability results established in Sections 2.3 and 4. These include polynomial growth potentials (Section 5.1), heavy-tailed marginals on general normed spaces (Section 5.2), boundary state-space models (Section 5.3), semi-compact and semi-discrete settings (Section 5.7), as well as finite mixtures of statistical models (Section 5.9). Section 5.8 is devoted to linear Gaussian models, where the regularity conditions and Lyapunov criteria admit explicit characterizations.
When the transition potential $W$ is uniformly bounded, weighted total variation norms do not provide additional insight. In this regime, exponential convergence of Sinkhorn semigroups is well understood through Hilbert’s projective metric and related contraction techniques [borwein, chen, deligiannidis, franklin], which already imply exponential decay in total variation.
On bounded subsets $\mathbb{X}\subset\mathbb{R}^{d}$, total variation also dominates Wasserstein distances associated with the Euclidean metric. However, the state space $\mathbb{X}$ is often non-complete and therefore non-Polish with respect to the Euclidean distance, which complicates Wasserstein convergence analysis. In this context, it is desirable to find a complete distance on the target space that makes the corresponding Wasserstein distance in distribution space complete [bolley2008separability].
This issue naturally motivates the introduction of alternative complete metrics. As a simple illustration, consider the Beta marginal distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
@@
Applying (2.10) to the total variation distance (see (1.14)) we readily check the Wasserstein exponential decays
@@
Although the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric, it becomes complete under
@@
for some $c_{\delta}>0$. In this context, the l.h.s. estimate (2.15) yields the $\varrho_{c,\delta}$-Wasserstein estimates
@@
Theorem 2.3 also applies to Gaussian marginal distributions on $\mathbb{Y}=\mathbb{X}=\mathbb{R}^{d}$ and a transition potential $W$ with at most quadratic growth (cf. Section 5.1). In this context, for $(p,q)=(2,1)$ in (1.11), we obtain exponential convergence in the the $2$-Wasserstein distance,
@@
Section 5.8 provides an explicit description of the functions $({\mathcal{U}}_{\delta},{\mathcal{V}}_{\delta})$ in the context of linear-Gaussian models. For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\mathcal{Q}$} with variance $\tau>2$ we have
${2}/{\tau}<\delta<1\Longrightarrow{\mathcal{U}}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$. }$
A symmetric Gaussian reference transition $\mathcal{Q}$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$. In this context, the exponential stability of Sinkhorn bridges w.r.t. the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [chiarini, durmus, del2025stability]. The article [adm-24] also provides Wasserstein exponential decays for any Gaussian marginals and any linear-Gaussian reference transition. For these classes of log-concave-type models, our Lyapunov approach yields new weighted total variation norm estimates for sufficiently large time horizons.
@@
Sections 2.3 and 5 demonstrate that Lyapunov techniques provide a flexible tool for establishing quantitative stability of Sinkhorn semigroups. Nevertheless, several important limitations remain.
@@
$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
This rather elementary example is a a slight modification of (2.12) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
@@
In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1.1). Nevertheless for uniformly bounded transition potential functions, Theorem 2.2 applies to any class of marginal distributions.
@@
3. Contraction of Markov operators